### 06 Mar 2024 v3

# atomic orbital API Application Programmer Interface

AO
atomic orbital
API
Application Programmer Interface
AUS
Advanced User Support
BEM
Boundary Element Method
BO
Born-Oppenheimer
CBS
complete basis set
CC
Coupled Cluster
CTCC
Centre for Theoretical and Computational Chemistry
CoE
Centre of Excellence
DC
dielectric continuum
DCHF
Dirac-Coulomb Hartree-Fock
DFT
density functional theory
DKH
Douglas-Kroll-Hess
EFP
effective fragment potential
ECP
effective core potential
EU
European Union
GGA
generalized gradient approximation
GPE
Generalized Poisson Equation
GTO
Gaussian Type Orbital
HF
Hartree-Fock
HPC
high-performance computing
HC
Hylleraas Centre for Quantum Molecular Sciences
IEF
Integral Equation Formalism
IGLO
individual gauge for localized orbitals
KB
kinetic balance
KS
Kohn-Sham
LAO
London atomic orbital
LAPW
linearized augmented plane wave
LDA
local density approximation
MAD
mean absolute deviation
maxAD
maximum absolute deviation
MM
molecular mechanics
MCSCF
multiconfiguration self consistent field
MPA
multiphoton absorption
MRA
multiresolution analysis
MSDD
Minnesota Solvent Descriptor Database
MW
multiwavelet
NAO
numerical atomic orbital
NeIC
nordic e-infrastructure collaboration
KAIN
Krylov-accelerated inexact Newton
NMR
nuclear magnetic resonance
NP
nanoparticle
NS
non-standard
OLED
organic light emitting diode
PAW
projector augmented wave
PBC
Periodic Boundary Condition
PCM
polarizable continuum model
PW
plane wave
QC
quantum chemistry
QM/MM
quantum mechanics/molecular mechanics
QM
quantum mechanics
RCN
Research Council of Norway
RMSD
root mean square deviation
RKB
restricted kinetic balance
SC
semiconductor
SCF
self-consistent field
STSM
short-term scientific mission
SAPT
symmetry-adapted perturbation theory
SERS
surface-enhanced raman scattering
WP1
Work Package 1
WP2
Work Package 2
WP3
Work Package 3
WP
Work Package
X2C
exact two-component
ZORA
zero-order relativistic approximation
ae
almost everywhere
BVP
boundary value problem
PDE
partial differential equation
RDM
1-body reduced density matrix
SCRF
self-consistent reaction field
IEFPCM
Integral Equation Formalism polarizable continuum model (PCM)
FMM
fast multipole method
DD
domain decomposition
TRS
time-reversal symmetry
SI
Supporting Information
DHF
Dirac–Hartree–Fock
MO
molecular orbital
[1,2]Christian Tantardini [1]Luca Frediani
1]Hylleraas center, Department of Chemistry, UiT The Arctic University of Norway, PO Box 6050 Langnes, N-9037 Tromsø, Norway. 2]Department of Materials Science and NanoEngineering, Rice University, Houston, Texas 77005, United States of America. 3]Algorithmiq Ltd., Kanavakatu 3C, FI-00160, Helsinki, Finland.
4-component Relativistic Calculations in a Multiwavelet Basis with Improved Convergence
christiantantardini@ymail.com    Roberto Di Remigio Eikås    luca.frediani@uit.no [ [ [
Abstract
The many-body wave function of an $N$-electron system within a relativistic framework can be described by the Dirac equation. Unfortunately, the Dirac operator $\hat{\mathfrak{D}}$ is unbounded. Thus, anions show the variational collapse of $N+1$ electron wavefunction making impossible to extend the $\hat{\mathfrak{D}}$ operator to such systems. Thus, it is necessary to avoid it and an alternative approach is based on applying the square of the Dirac operator $\hat{\mathfrak{D}}^{2}$. This approach is especially suitable for a multiwavelet framework: its implementation in an integral equation form is readily available and the completeness of the multiwavelet basis affords tight error control. The use of the $\hat{\mathfrak{D}}^{2}$ operator within a multiwavelet (MW) framework is able to achieve higher precision compared to the $\hat{\mathfrak{D}}$, consistent with the quadratic nature of the error estimates in a variational framework.
keywords:
, multiwavelets, Dirac equation, variational collapse, squared-Dirac operator
1 Introduction
Dirac’s equation describes the relativistic motion of electrons and still today represents the most accurate physical model, short of quantum electrodynamics, for the inclusion of relativistic effects in molecular calculations. Paul Dirac formulated his equation by aiming for a quantum mechanical equation that satisfied Lorentz covariance, while guaranteeing a positive semi-definite probability density.[1]
In its original formulation, the Dirac equation suffers from some drawbacks and the most important one is called variational collapse, a term coined by Schwarz and Wallmeier [2], because the Dirac operator $\hat{\mathfrak{D}}$, in contrast to the Schrödinger Hamiltonian $\hat{H}$, is not bounded from below giving spurious unphysical solutions as the kinetic energy tends to zero. Therefore, the application of variational methods is not justified. Schwarz and Wallmeier [2] were the first to observe: (i) the collapsing of two neighboring exponents of basis sets to the same value during optimization; (ii) no systematic variation of the eigenvalues with increasing basis size (n.b. eigenvalues change with basis set). During the years the same was observed by others [3, 4, 5], and recently with contracted Dyall basis sets [6, 7, 8, 9, 10] and atomic-balanced basis sets [11, 12, 13].
Such effects were attenuated by expanding the basis sets through a Legendre parameterization which provides an efficient way of systematically approaching the basis set limit and by using the restricted kinetic balance (RKB) to generate the small components of the spinorbitals [14, 15, 16]. However, the presence of uncontracted basis sets with the RKB could generate the prolapse due to the lack of kinetic balance in the description of the negative-energy states [17, 14, 11, 18, 19].
Nowadays, the variational collapse can be considered largely mitigated by the use of appropriate atomic basis sets in conjunction with the RKB prescription, but the problem continues to exist for electronic metastable anions, where the four-component Dirac wavefunction of $N+1$ electron will collapse in an orbital that is infinitely removed from the parent plus forming a continuum function [20].
There are essentially four strategies to deal with the variational collapse:
  1. 1.
    accepting variational collapse and trying to minimize its effects by suppressing the small components and introducing $c\rightarrow\infty$ [3, 21, 22] or special choice of basis set [23];
  2. 2.
    transforming the Dirac operator (e.g., Foldy-Wouthuysen (FW) transformation [24, 25]) to achieve a matrix representation that incorporates the correct non-relativistic limit;
  3. 3.
    manipulating the matrix representation of the Dirac operator (e.g., by modifying $\vec{\sigma}\cdot\vec{p}$ [5], by the stabilization method [26, 27] or the rotation method [28]) without transforming the operator itself;
  4. 4.
    replacing the Dirac operator by a related operator which has the same eigenfunctions, but which is bounded from below such that the variation principle can be applied (e.g., squared Dirac operator [4] or projection of the Dirac operator to positive energy states [29, 30]).
Here, we will pursue the latter option in the field of adaptive multiwavelet basis. The use of $\hat{\mathfrak{D}}^{2}$ was originally presented by Wallmeier and Kutzelnigg [4] for one-electron calculations and later extended to molecular calculations [31]. We first review the theory of $\hat{\mathfrak{D}}^{2}$ for one- and many-electron systems. For the latter, we introduce the unretarded Coulombic interaction to describe the repulsion between electrons, leading to the direct ($\hat{J}$) and exchange ($\hat{K}$) interactions in the mean-field case. We then apply this method to simple one- and two-electron systems to show the viability of the approach and the precision which can be attained compared to the standard Dirac equation.
2 Theoretical Background
In the following, we will use Hartree atomic units. However, for clarity, we will keep the electron mass $m$ explicit in the equations. The free-space Dirac operator is given as:
$\hat{h}_{D}=c\vec{\alpha}\cdot\vec{p}+\beta mc^{2}$ (1)
where $\vec{p}$ collects the Cartesian components of the momentum operator:
$\vec{p}=-\mathrm{i}\begin{pmatrix}\frac{\partial}{\partial x}\\ \frac{\partial}{\partial y}\\ \frac{\partial}{\partial z}\end{pmatrix}$ (2)
and $\vec{\alpha}$ is a Cartesian vector whose components are $4\times 4$ anti-diagonal block matrices:
$\alpha_{u}=\begin{pmatrix}0&\sigma_{u}\\ \sigma_{u}&0\end{pmatrix}$ (3)
We use indices $u,w\in\{x,y,z\}$ for Cartesian components. The $2\times 2$ matrices $\sigma_{u},\,u\in\{x,y,z\}$ appearing as anti-diagonal blocks are the Pauli matrices. The $4\times 4$ matrix $\beta$ is defined as:
$\beta=\begin{pmatrix}I_{2}&0\\ 0&-I_{2}\end{pmatrix}$ (4)
The eigenfunctions of the free-space Dirac operator are vectors with four complex components:
$\Phi=\begin{pmatrix}\varphi^{1}\\ \varphi^{2}\\ \varphi^{3}\\ \varphi^{4}\end{pmatrix},$ (5)
the upper (lower) two components of the orbital can equivalently be referred to as large (small) components and indexed accordingly as:
$\Phi=\begin{pmatrix}\varphi^{L_{\alpha}}\\ \varphi^{L_{\beta}}\\ \varphi^{S_{\alpha}}\\ \varphi^{S_{\beta}}\end{pmatrix}.$ (6)
The large and small components can be transformed into each other. Practically, the exact relationship is only applicable for the free-particle problem. In atomic and molecular applications, one initializes the small components by applying the kinetic balance condition, i.e. the nonrelativistic limit of the exact large-small relationship, to the large components [32]. In the following, the four-component molecular orbitals will further be indexed with a subscript Latin letters, e.g. $p,q,i,j,\ldots\in\{1,\ldots,N\}$, with $N$ the number of such orbitals.
2.1 Adaptive Multiwavelets Basis
[…]
MRA is a numerical framework devised to solve integral and differential equations with arbitrary, predefined precision.[33] Standard quantum chemical methods expand the unknown molecular orbitals into fixed atom-centered bases and recast the solution of the Hartree–Fock equations as a matrix eigenvalue problem to determine the expansion coefficients [34]. multiresolution analysis (MRA)-based methods instead solve directly for the occupied MOs, through adaptive refinement of the expansion basis. The mathematical framework enabling such adaptive algorithms relies on the construction of a sequence of orthogonal subspaces which is dense in the space of square-integrable functions for a user-defined simulation box.
As an example, consider the unit interval and a set of orthonormal $k$-th order polynomials, $\phi_{i}(x)$, which we call scaling functions. The basis can be refined by dyadic subdivisions of the interval: at any subdivision scale, $n$, the original unit interval is subdivided into $2^{n}$ sub-intervals. Each interval supports a fixed number of scaling functions, which are obtained by dilation and translation of the same functions at scale $n=0$:
$\phi^{n}_{il}=2^{n/2}\phi_{i}(2^{n}x-l),$ (7)
where $l=0,\ldots,2^{n-1}$ is the translation index, identifying functions in the sub-interval $[l/2^{n},(l+1)/2^{n}]$. This construction defines a telescopic series of scaling spaces:
$V_{0}^{k}\subset V_{1}^{k}\subset.....V_{n}^{k}\subset....\subset L^{2}([0,1]),$ (8)
The limit of this series is dense in $L^{2}([0,1])$ and can thus represent any $f\in L^{2}([0,1])$. The orthogonal complement between two successive scaling spaces constitutes a wavelet space:
$S_{n}^{k}\oplus W_{n}^{k}=S_{n+1}^{k},\>\>\>W_{n}^{k}\perp S_{n}^{k},$ (9)
The crucial property of wavelet functions is constituted by their vanishing moments, which are a straightforward consequence of their construction as orthogonal complement of a set of polynomial functions. This is the key feature that leads to sparse representations of functions, enabling both precision control on the one hand and fast algorithms on the other. The $k$ parameter is the order of the Legendre polynomials used in each refinement box, which is regardless of the number of of electrons in the system and, in principle, a given a user-defined tolerance, $\epsilon$, which is used both as requested precision in the adaptive refinement and operator application algorithms, and as convergence tolerance of the self-consistent field (SCF) cycle. Convergence is achieved once the norm of the orbital update drops below $\epsilon$. This means that if the error on the orbital is $10^{-N}$ the error on the energy will be $10^{-2N}$.
We note that the MRA construction is dense in $L^{2}$ not because of the high polynomial order, but because of the telescopic construction of subspaces by dyadic subdivision. In principle, one could achieve completeness of the representation basis even using $k=1$, the telescopic series of spaces would however be impractically large. In practice, as we showed in our previous work on the full Breit Hamiltonian [35], requesting tighter precision in a relativistic setting demands a higher polynomial order than what provided by default by MRCPP, our MRA library. The interested reader is referred to the available literature for details about MRA and its practical realization for quantum chemistry applications.[36, 37, 38, 39, 40]
2.2 Integral equation formulation for the Dirac operator
The Dirac-Fock equations are the 4-component, relativistic counterpart of the Hartree-Fock equations:
$\hat{\mathfrak{D}}\Phi_{i}=(\hat{h}_{D}+\hat{V})\Phi_{i}=(c\vec{\alpha}\cdot\vec{p}+\beta mc^{2}+\hat{V})\Phi_{i}=\sum_{j}D_{ij}\Phi_{j},$ (10)
where $\hat{h}_{D}=c\vec{\alpha}\cdot\vec{p}+\beta mc^{2}$ is the free-particle Dirac operator and $\hat{V}=\hat{J}-\hat{K}+\hat{V}_{\mathrm{nuc}}$ is the SCF one-body potential. The matrix $D_{ij}=\braket{\Phi_{i}}{\hat{\mathfrak{D}}}{\Phi_{j}}$ is the projection of the mean-field Hamiltonian $\hat{\mathfrak{D}}$ in the occupied subspace spanned by $N$ orthonormal spinors. These are a set of coupled first-order differential equations. Its solutions are 4-component spinors; each component is a complex function and the set of solutions describes both positive and negative energy states. We rearrange Eq. (10) as:
$(\hat{h}_{D}-D_{ii})\Phi_{i}=-\hat{V}\Phi_{i}+\sum_{j\neq i}D_{ij}\Phi_{j}.$ (11)
The nonrelativistic Schrödinger equation, and its mean-field equivalents, can be solved by iterative convolution with the Green’s function for the shifted Laplacian [41]. Following Blackledge and Babajanov [42], we formulate a similar iterative convolution for Eq. (11). First, we introduce the identity on the left-hand side:
$(\hat{h}_{D}-D_{ii})(\hat{h}_{D}+D_{ii})(\hat{h}_{D}+D_{ii})^{-1}\Phi_{i}$ $=\left[\hat{h}_{D}^{2}-D_{ii}^{2}\right](\hat{h}_{D}+D_{ii})^{-1}\Phi_{i}$ (12)
$=-\hat{V}\phi_{i}+\sum_{j\neq i}D_{ij}\Phi_{j},$
The term in square brackets, thanks to Dirac’s identity $(\vec{\sigma}\cdot\vec{p})(\vec{\sigma}\cdot\vec{p})=p^{2}$, is diagonal in the four components:
$\hat{h}_{D}^{2}-D_{ii}^{2}=\left[-c^{2}\nabla^{2}+m^{2}c^{4}-D_{ii}^{2}\right]I_{4}.$ (13)
This yields:
$\left[-\nabla^{2}+\frac{m^{2}c^{4}-D_{ii}^{2}}{c^{2}}\right]I_{4}(\hat{h}_{D}+D_{ii})^{-1}\Phi_{i}=-\frac{1}{c^{2}}\hat{V}\Phi_{i}+\frac{1}{c^{2}}\sum_{j\neq i}D_{ij}\Phi_{j},$ (14)
which can be formally inverted to give:[43]
$\Phi_{i}=-\frac{1}{c^{2}}\left[(\hat{h}_{D}+D_{ii})G_{i}\right]\star\left[\hat{V}\Phi_{i}-\sum_{j\neq i}D_{ij}\Phi_{j}\right],$ (15)
where $G_{i}$ is the Helmholtz kernel:
$G_{i}\equiv\left[-\nabla^{2}+\frac{m^{2}c^{4}-D_{ii}^{2}}{c^{2}}\right]^{-1}I_{4}=I_{4}\frac{1}{4\pi}\frac{\exp{(-\mu_{i}|\vec{r}-\vec{r}^{\prime}|)}}{|\vec{r}-\vec{r}^{\prime}|}$ (16)
with exponent:
$\mu_{i}=\sqrt{\frac{m^{2}c^{4}-D_{ii}^{2}}{c^{2}}}>0,$ (17)
since we are seeking bound states, i.e. $D_{ii}<mc^{2}$. As noted by Anderson et al. [40], one can use the properties of the convolution product to rearrange Eq. (15) as:
$\left[(\hat{h}_{D}+D_{ii})G_{i}\right]\star\left[\hat{V}\Phi_{i}-\sum_{j\neq i}D_{ij}\Phi_{j}\right]$ (18a)
$(\hat{h}_{D}+D_{ii})\left[G_{i}\star\left(\hat{V}\Phi_{i}-\sum_{j\neq i}D_{ij}\Phi_{j}\right)\right]$ (18b)
$G_{i}\star\left[(\hat{h}_{D}+D_{ii})\left(\hat{V}\Phi_{i}-\sum_{j\neq i}D_{ij}\Phi_{j}\right)\right]$ (18c)
See also Appendix A for a detailed derivation of these equivalences. As we show in Appendix B, the iterative convolution of Eq. (18a) uses the inverse-distance-cube kernel, which also appears in the context of the relativistic two-electron gauge potential [44]. As shown by us in Ref. [35], application of this kernel is computationally intensive and not numerically robust. The second form, Eq. (18b), is computationally more efficient: we apply the same convolution kernel used in the nonrelativistic case first, smoothing out any sharp features introduced by multiplication with the potential, and then we compute its derivatives [45]. The final form in Eq. (18c) requires application of the derivative operator on $V\phi_{i}$, which might have non-smooth features and thus introduce numerical errors, convergence slowdown or excessive memory demands.
3 Developing of Squared Dirac Operator within Multiwavelets Field
Following Wallmeier and Kutzelnigg [4], the square of the Dirac operator is positive and bounded from below, while obviously sharing eigenfunctions and eigenvalues with the original operator:
$\hat{\mathfrak{D}}^{2}\Phi_{i}$ =$=\left\{c^{2}p^{2}I_{4}+c[\vec{\alpha}\cdot\vec{p},\hat{V}]_{+}+m^{2}c^{4}+mc^{2}[\beta,\hat{V}]_{+}+\hat{V}^{2}I_{4}\right\}\Phi_{i}$ (19)
$=\sum_{j}(\hat{\mathfrak{D}}^{2})_{ij}\Phi_{j}.$
Shifting by $m^{2}c^{4}$ and rescaling by $2mc^{2}$ yields:
$\left\{\frac{p^{2}}{2m}I_{4}+\hat{\mathcal{V}}\right\}\Phi_{i}=\frac{1}{2mc^{2}}\sum_{j}\left[(\hat{\mathfrak{D}}^{2})_{ij}-\delta_{ij}m^{2}c^{4}\right]\Phi_{j},$ (20)
where, for brevity, we define the generalized mean-field potential:
$\hat{\mathcal{V}}=\frac{[\beta,\hat{V}]_{+}}{2}+\frac{[\vec{\alpha}\cdot\vec{p},\hat{V}]_{+}}{2mc}+\frac{\hat{V}^{2}}{2mc^{2}}I_{4}.$ (21)
At convergence, the positive-energy orbital energies of $\hat{\mathfrak{D}}$ are $\epsilon_{i}=-mc^{2}+\sqrt{m^{2}c^{4}+2mc^{2}\omega_{i}}$, where $\omega_{i}$ are the eigenvalues of the shifted-and-scaled SCF equations above. Eq. (20) is close to the nonrelativistic Hartree-Fock (HF) equations, despite the appearance of more complicated potential-like terms. Similarly to Eq. (10) it can be recast as an integral equation and solved by iterative convolution:
$\Phi_{i}=-2mG_{i}\star\left[\mathcal{V}\Phi_{i}+\frac{1}{2mc^{2}}\sum_{j\neq i}(\hat{\mathfrak{D}}^{2})_{ij}\Phi_{j}\right],$ (22)
We choose to approximate the matrix elements of the $\hat{\mathfrak{D}}^{2}$ operator as:
$(\hat{\mathfrak{D}}^{2})_{ij}\simeq\sum_{k}D_{ik}D_{kj},$ (23)
in the converged occupied subspace and for a complete basis (MWs are to be regarded as a complete basis to within their predefined precision), this relation would be exact.111 The MRA construction only explicitly refines the occupied orbital space and the SCF procedure iteratively block-diagonalizes the matrix representation of the Fock operator. Away from convergence, the off-diagonal blocks will be nonzero: $\begin{pmatrix}\mathbf{O}&\mathbf{A}\\ \mathbf{A}^{\dagger}&\mathbf{V}\end{pmatrix}.$ Forming the square of the Fock operator as the product gives: $\begin{pmatrix}\mathbf{O}^{2}+\mathbf{A}\mathbf{A}^{\dagger}&\mathbf{O}\mathbf{A}+\mathbf{A}\mathbf{V}\\ \mathbf{A}^{\dagger}\mathbf{O}+\mathbf{V}\mathbf{A}^{\dagger}&\mathbf{V}^{2}+\mathbf{A}^{\dagger}\mathbf{A}\end{pmatrix}$ showing that only when $\mathbf{A}=0$ (at convergence) the representation in the occupied basis of $\mathbf{F}^{2}$ will be equal to the product of the basis representation of $\mathbf{F}$ with itself. The exponent $\mu_{i}$ in Eq. (17) is now:
$\mu_{i}=\sqrt{\frac{m^{2}c^{4}-(D^{2})_{ii}}{c^{2}}}\simeq\sqrt{\frac{m^{2}c^{4}-\sum_{k}D_{ik}D_{ki}}{c^{2}}}.$ (24)
For the positive-energy states of interest, it always holds that $\mu_{i}>0$.
3.1 The generalized mean-field potential
The iterative solution of the integral equation (22) requires the repeated application of the generalized mean-field potential in Eq. (21). For general operators in a spinor basis it holds:
$[\beta,M]_{+}=2\begin{pmatrix}M^{L_{\alpha}L_{\alpha}}&M^{L_{\alpha}L_{\beta}}&0&0\\ M^{L_{\beta}L_{\alpha}}&M^{L_{\beta}L_{\beta}}&0&0\\ 0&0&-M^{S_{\alpha}S_{\alpha}}&-M^{S_{\alpha}S_{\beta}}\\ 0&0&-M^{S_{\beta}S_{\alpha}}&-M^{S_{\beta}S_{\beta}}\end{pmatrix}.$ (25)
In a MW framework, we work directly with the MO basis: thus the potential operator $V$ is diagonal in the spinor components and its anticommutator with $\beta$ is simply:
$[\beta,\hat{V}]_{+}=2\beta\hat{V}.$ (26)
Application of $\hat{\mathcal{V}}$ involves the, potentially demanding, construction of the squared mean-field potential. Formally, its application on a test spinor is:
$\hat{V}^{2}\Phi_{k}=\hat{J}^{2}\Phi_{k}+\hat{K}^{2}\Phi_{k}+\hat{V}_{\mathrm{nuc}}^{2}\Phi_{k}+[\hat{J}-\hat{K},\hat{V}_{\mathrm{nuc}}]_{+}\Phi_{k}-[\hat{J},\hat{K}]_{+}\Phi_{k},$ (27)
however, since we are working in a first-quantization setting, these operators can be applied as:
$\hat{V}^{2}\Phi_{k}=\hat{V}\left(\hat{V}\Phi_{k}\right).$ (28)
The three terms in Eq. (21) enter at different orders in $\frac{1}{c}$. A possible performance gain, especially when working with larger molecular systems, would be to apply them with a threshold obtained by scaling the user-defined one with their respective denominators.
For the one- and two-electron systems we consider in this work, the generalized potential is greatly simplified. For one-electron systems, we have:
$\hat{\mathcal{V}}=\beta\hat{V}_{\mathrm{nuc}}+\frac{[\vec{\alpha}\cdot\vec{p},\hat{V}_{\mathrm{nuc}}]_{+}}{2mc}+\frac{\hat{V}_{\mathrm{nuc}}^{2}}{2mc^{2}}I_{4},$ (29)
For closed-shell, two-electron systems, we can explicitly exploit Kramers’ time-reversal symmetry (TRS) and remove the exchange terms:[46]
$\hat{\mathcal{V}}=\beta(\hat{J}+\hat{V}_{\mathrm{nuc}})+\frac{[\vec{\alpha}\cdot\vec{p},\hat{J}+\hat{V}_{\mathrm{nuc}}]_{+}}{2mc}+\frac{(\hat{J}+\hat{V}_{\mathrm{nuc}})^{2}}{2mc^{2}}I_{4},$ (30)
4 Results and Discussion
All MW results presented were obtained with our Python implementation in the ReMRChem project, whose source code is freely avaialble through GitHub. [^2] https://github.com/MRChemSoft/ReMRChem. Our implementation uses the MRCPP library [47] and its Python interface provided by the VAMPyR package [48, 49]. We have set up our calculations choosing the smeared nuclear point-like charge, which replaces the nuclear $1/r$ potential with a smoothed approximation[50]. This smoothing has no physical meaning and it must not be misunderstood with the physical models introduced by Visscher [51], which are common in relativistic methods. The value of the speed of light was chosen equal to 137.03599913900001 a.u. and for the calculation of the derivative we used the B-spline method introduced by Anderson et al.[52].
The current implementation uses the VAMPyR package (Python code) and is a proof-of-concept implementation to validate the use of the squared Dirac operator within the multiwavelets framework. This implementation can only deal with at most two-electron systems since it relies, as already stated, on Kramers’ TRS [46, 53] to reduce memory footprint and computational load. This is the ideal starting point to consider inter-electronic interactions in the SCF setting for the validation of the developed theory. However, as it is known from the nonrelativistic MW implementation in the MRChem package (written in C++), increasing the number of electrons will not pose additional theoretical challenges, but only increase the required computational resources (memory in particular) as tested by our group and extensively discussed in Ref. 54.
[…]
Our results show that at low precision the two implementations are equivalent, when precision is increased, the $\hat{\mathfrak{D}}^{2}$ implementation leads to slightly lower errors (up to one order of magnitude) than the $\hat{\mathfrak{D}}$ implementation.
[…]
$\hat{\mathfrak{D}}(\epsilon=10^{-4})$0.500007467791874 $|1.62236|\cdot 10^{-06}$1.102640862147382 $|6.51808|\cdot 10^{-07}$
$\hat{\mathfrak{D}}^{2}(\epsilon=10^{-4})$0.500007785136819 $|2.25705|\cdot 10^{-06}$1.102640727757642 $|7.73929|\cdot 10^{-07}$
$\hat{\mathfrak{D}}(\epsilon=10^{-6})$0.500006653182936 $|6.83203|\cdot 10^{-09}$1.102641514193238 $|6.06142|\cdot 10^{-08}$
$\hat{\mathfrak{D}}^{2}(\epsilon=10^{-6})$0.500006655675723 $|1.84653|\cdot 10^{-09}$1.102641581857073 $|7.47750|\cdot 10^{-10}$
$\hat{\mathfrak{D}}(\epsilon=10^{-8})$0.500006656217010 $|7.63966|\cdot 10^{-10}$1.102641581459466 $|3.77258|\cdot 10^{-10}$
$\hat{\mathfrak{D}}^{2}(\epsilon=10^{-8})$0.500006656559709 $|7.85770|\cdot 10^{-11}$1.102641580981426 $|4.63886|\cdot 10^{-11}$
Exact result -0.500006656598998 -1.102641581032577
[…]
22footnotetext: Legendre polynomial order $k=10$ and tolerance $\epsilon=10^{-8}$, Legendre polynomial order $k=8$ and tolerance $\epsilon=10^{-6}$, Legendre polynomial order $k=6$ and tolerance $\epsilon=10^{-4}$.
[…]
We report the comparison of our results with radial numerical results previously obtained using GRASP Ref. 35 in Tab. 2. Also in this case we have performed calculations by increasing requested precision both for the $\hat{\mathfrak{D}}$ and $\hat{\mathfrak{D}}^{2}$ formulations. We observe that both implementations are able to achieve the requested precision. However, there is no noticeable difference between them.
[…]
Operator orbital energy Abs Rel. Error Total energy Abs Rel. Error
VAMPyR / (Ha) VAMPyR / (Ha)
$\hat{\mathfrak{D}}(\epsilon=10^{-4})$0.918040748 $|5.45383|\cdot 10^{-05}$2.861854359 $|1.43330|\cdot 10^{-05}$
$\hat{\mathfrak{D}}^{2}(\epsilon=10^{-4})$0.918032833 $|4.59170|\cdot 10^{-05}$2.861828187 $|5.18788|\cdot 10^{-06}$
$\hat{\mathfrak{D}}(\epsilon=10^{-6})$0.918008207 $|1.90908|\cdot 10^{-05}$2.861833848 $|7.16586|\cdot 10^{-06}$
$\hat{\mathfrak{D}}^{2}(\epsilon=10^{-6})$0.918008259 $|1.91475|\cdot 10^{-05}$2.861833976 $|7.21072|\cdot 10^{-06}$
$\hat{\mathfrak{D}}(\epsilon=10^{-8})$0.917990690 $|8.78784|\cdot 10^{-09}$2.861813340 $|1.93716|\cdot 10^{-10}$
$\hat{\mathfrak{D}}^{2}(\epsilon=10^{-8})$0.917990691 $|9.69563|\cdot 10^{-09}$2.861813342 $|5.19185|\cdot 10^{-10}$
GRASP Reference -0.917990682 -2.861813340
[…]
22footnotetext: Legendre polynomial order $k=10$ and tolerance $\epsilon=10^{-8}$, Legendre polynomial order $k=8$ and tolerance $\epsilon=10^{-6}$, Legendre polynomial order $k=6$ and tolerance $\epsilon=10^{-4}$.
[…]
Operator orbital energy Abs Rel. Error Total energy Abs Rel. Error
VAMPyR / (Ha) VAMPyR / (Ha)
$\hat{\mathfrak{D}}(\epsilon=10^{-4})$44.094786197 $|2.794189|\cdot 10^{-03}$94.108543063 $|1.337941|\cdot 10^{-03}$
$\hat{\mathfrak{D}}^{2}(\epsilon=10^{-4})$43.978792210 $|1.562788|\cdot 10^{-04}$93.990113164 $|7.781809|\cdot 10^{-05}$
$\hat{\mathfrak{D}}(\epsilon=10^{-6})$43.976907059 $|1.134071|\cdot 10^{-04}$93.987884382 $|5.410330|\cdot 10^{-05}$
$\hat{\mathfrak{D}}^{2}(\epsilon=10^{-6})$43.971930116 $|2.224998|\cdot 10^{-07}$93.982819929 $|2.162733|\cdot 10^{-07}$
$\hat{\mathfrak{D}}(\epsilon=10^{-8})$43.971921492 $|2.638859|\cdot 10^{-08}$93.982804435 $|5.142152|\cdot 10^{-08}$
$\hat{\mathfrak{D}}^{2}(\epsilon=10^{-8})$43.971921513 $|2.686178|\cdot 10^{-08}$93.982804554 $|5.268882|\cdot 10^{-08}$
GRASP Reference -43.971920332 -93.982799603
[…]
22footnotetext: Legendre polynomial order $k=10$ and tolerance $\epsilon=10^{-8}$, Legendre polynomial order $k=8$ and tolerance $\epsilon=10^{-6}$, Legendre polynomial order $k=6$ and tolerance $\epsilon=10^{-4}$.
[…]
5 Conclusions
The implementation of the squared Dirac operator $\hat{\mathfrak{D}}^{2}$ within a MW framework allows the use of the same integral equation formalism which has been widely tested in the context of non-relativistic theory. The original papers reporting the use of such operators [4, 31], indicated the importance of making use of a large (by practical means complete) basis set to achieve valid results with this method. The main reason is to make sure that the matrix representation of the squared operator is equivalent to the matrix product in Eq. (23). Such a requirement was challenging to satisfy in an atom-centered basis with the limited computational resources available in the early 1980s, and to some extent it is not trivial to achieve for Gaussian Type Orbitals. On the other hand, MWs have by and large demonstrated that they can be regarded as a complete basis set.
In the present work, we have shown that the use of the $\hat{\mathfrak{D}}^{2}$ operator with MWs is possible and it provides results with an error with respect to radial numerical approach (i.e., GRASP) of the same order of magnitude as using the $\hat{\mathfrak{D}}$ operator.
Acknowledgments
We would like to thank Prof. Dr. Trond Saue from the CNRS/Université de Toulouse, France. We acknowledge support from the Research Council of Norway through its Centres of Excellence scheme (262695), through the FRIPRO grant ReMRChem (324590), and from NOTUR – The Norwegian Metacenter for Computational Science through grant of computer time (nn4654k).
Declarations
We use the CRediT taxonomy of contributor roles.[57, 58] The “Investigation” role also includes the “Methodology”, “Software”, and “Validation” roles. The “Analysis” role also includes the “Formal analysis” and “Visualization” roles. The “Funding acquisition” role also includes the “Resources” role. We visualize contributor roles in the following authorship attribution matrix, as suggested in Ref. 59.
Table 4: Levels of contribution: major, minor.
CT RDRE LF
Conceptualization
Investigation
Data curation
Analysis
Supervision
Writing – original draft
Writing – revisions
Funding acquisition
Project administration
Appendix A Equivalence of integral operator iterations
The equivalence of the integral operator iterations in Eqs. (18a), (18b), and (18c) rests on the relation between differentiation and convolution products:
$(f\star g)^{\prime}=f\star g^{\prime}.$ (31)
We use square brackets to delimit the operands in the convolution. Expanding Eq. (18a):
$\Phi_{i}=$ $-\frac{1}{c^{2}}\left[c(\vec{\alpha}\cdot\vec{p})G_{i}\right]\star\left[V\Phi_{i}-\sum_{j\neq i}D_{ij}\Phi_{j}\right]$ (32)
$-\frac{1}{c^{2}}\beta mc^{2}\left[G_{i}\right]\star\left[V\Phi_{i}-\sum_{j\neq i}D_{ij}\Phi_{j}\right]$
$-\frac{D_{ii}}{c^{2}}\left[G_{i}\right]\star\left[V\Phi_{i}-\sum_{j\neq i}D_{ij}\Phi_{j}\right],$
where both the scalar and $\beta mc^{2}$ terms have been moved outside of the convolution, i.e. outside of the square brackets, the latter since $\beta I_{4}=I_{4}\beta$. The first term in the convolution contains the derivative of $G_{i}$, thus we can use property (31) to move it outside of the convolution:
$\Phi_{i}=$ $-\frac{1}{c^{2}}c(\vec{\alpha}\cdot\vec{p})\left[G_{i}\right]\star\left[V\Phi_{i}-\sum_{j\neq i}D_{ij}\Phi_{j}\right]$ (33)
$-\frac{1}{c^{2}}\beta mc^{2}\left[G_{i}\right]\star\left[V\Phi_{i}-\sum_{j\neq i}D_{ij}\Phi_{j}\right]$
$-\frac{D_{ii}}{c^{2}}\left[G_{i}\right]\star\left[V\Phi_{i}-\sum_{j\neq i}D_{ij}\Phi_{j}\right],$
obtaining Eq. (18b), after rearrangement. From this form, we can use property (31) once more to move the derivative application on the $V\phi_{i}$ convolution partner,
$\Phi_{i}=$ $-\frac{1}{c^{2}}\left[G_{i}\right]\star\left[c(\vec{\alpha}\cdot\vec{p})\left(V\Phi_{i}-\sum_{j\neq i}D_{ij}\Phi_{j}\right)\right]$ (34)
$-\frac{1}{c^{2}}\beta mc^{2}\left[G_{i}\right]\star\left[V\Phi_{i}-\sum_{j\neq i}D_{ij}\Phi_{j}\right]$
$-\frac{D_{ii}}{c^{2}}\left[G_{i}\right]\star\left[V\Phi_{i}-\sum_{j\neq i}D_{ij}\Phi_{j}\right],$
Eq. (18c) follows, since the corresponding rearrangement for the additional two terms in the convolution is trivial.
Appendix B Explicit form of the convolution kernel in Eq. (16a)
Whereas the iterative integral operator application in Eqs. (18b) and (18c) uses the (diagonal) matrix form of the nonrelativistic bound-state Helmholtz kernel, the iteration in Eq. (18a) requires the use of a new convolution kernel: $c(\vec{\alpha}\cdot\vec{p})G_{i}$. This is a $4\times 4$ matrix convolution kernel:
$c(\vec{\alpha}\cdot\vec{p})G_{i}=c\begin{pmatrix}0&0&\partial_{z}g_{i}&(\partial_{x}-\mathrm{i}\partial_{y})g_{i}\\ 0&0&(\partial_{x}+\mathrm{i}\partial_{y})g_{i}&-\partial_{z}g_{i}\\ \partial_{z}g_{i}&(\partial_{x}-\mathrm{i}\partial_{y})g_{i}&0&0\\ (\partial_{x}+\mathrm{i}\partial_{y})g_{i}&-\partial_{z}g_{i}&0&0\end{pmatrix}$ (35)
with $g_{i}=\frac{1}{4\pi}\frac{\exp{(-\mu_{i}|\vec{r}-\vec{r}^{\prime}|)}}{|\vec{r}-\vec{r}^{\prime}|}$. The derivatives of the bound-state Helmholtz kernel are:
$\partial_{u}g_{i}$ $=\frac{1}{4\pi}\left(\mu_{i}\frac{(r_{u}-r_{u}^{\prime})\exp{(-\mu_{i}|\vec{r}-\vec{r}^{\prime}|)}}{|\vec{r}-\vec{r}^{\prime}|^{2}}+\frac{(r_{u}-r_{u}^{\prime})\exp{(-\mu_{i}|\vec{r}-\vec{r}^{\prime}|)}}{|\vec{r}-\vec{r}^{\prime}|^{3}}\right)$ (36)
$=\frac{1}{4\pi}\left[\left(\mu_{i}+\frac{1}{|\vec{r}-\vec{r}^{\prime}|}\right)\frac{1}{|\vec{r}-\vec{r}^{\prime}|^{2}}\right](r_{u}-r_{u}^{\prime})\exp{(-\mu_{i}|\vec{r}-\vec{r}^{\prime}|)},\quad u\in\{x,y,z\},$
leading to a $4\times 4$, anisotropic convolution kernel:
$c(\vec{\alpha}\cdot\vec{p})G_{i}$ $=\frac{c}{4\pi}\left[\left(\mu_{i}+\frac{1}{|\vec{r}-\vec{r}^{\prime}|}\right)\frac{1}{|\vec{r}-\vec{r}^{\prime}|^{2}}\right]\exp{(-\mu_{i}|\vec{r}-\vec{r}^{\prime}|)}\times$ (37)
$\begin{pmatrix}0&0&(r_{z}-r_{z}^{\prime})&(r_{x}-\mathrm{i}r_{y})-(r_{x}^{\prime}-\mathrm{i}r_{y}^{\prime})\\ 0&0&(r_{x}+\mathrm{i}r_{y})-(r_{x}^{\prime}+\mathrm{i}r_{y}^{\prime})&-(r_{z}-r_{z}^{\prime})\\ (r_{z}-r_{z}^{\prime})&(r_{x}-\mathrm{i}r_{y})-(r_{x}^{\prime}-\mathrm{i}r_{y}^{\prime})&0&0\\ (r_{x}+\mathrm{i}r_{y})-(r_{x}^{\prime}+\mathrm{i}r_{y}^{\prime})&-(r_{z}-r_{z}^{\prime})&0&0\end{pmatrix}$